OpenStudy (anonymous):

If measures 135, what is the length of the radius of this circle? Round your answer to two decimal places.

OpenStudy (phi):

I take it you did not understand the answer I posted to this question?

OpenStudy (anonymous):

@phi i decided to skip it but i failed the test

OpenStudy (phi):

oh. If you learn how to do these problems, you could pass. what part is confusing you ?

OpenStudy (anonymous):

i went back and read it just dont know how to solve it:(

OpenStudy (gudden):

use the formula... (theta/360)2 pie r = 9

OpenStudy (anonymous):

theta?

OpenStudy (phi):

Here is the main idea: a circle has a circumference (distance around it) a circle is 360 degrees. if you walk in a circle for 180º (for example) you only walked part way around the circle. because 180 is ½ of 360, that means you walked half way around. if you walked 120 degrees that is 120/360 = ⅓ of the way around the circle.

OpenStudy (phi):

so you should be able to figure out: if I go 90º around a circle, what fraction of the circle is that ?

OpenStudy (gudden):

thetha means an angle...which in this case is 135 degree

OpenStudy (anonymous):

1/4?

OpenStudy (phi):

yes, exactly. sudden posted the general formula. if we call the angle theta (90 in this case) the fraction around the circle is theta/360. 90/360 in this example= ¼

OpenStudy (phi):

*gudden

OpenStudy (phi):

so think about that idea for a bit: if you walk 30º out of 360º you walked 30/360 =3/36= 1/12 of the way (not very far)

OpenStudy (phi):

if you walk 360 degrees you walked around the circle exactly once.

OpenStudy (phi):

next, you probably should learn that the circumference (length around the circle) is C= 2 pi r where pi is 3.14159... (use 3.14 is a good choice) and r is the radius

OpenStudy (phi):

You really should practice on the earlier problems before tackling this one If measures 135, what is the length of the radius of this circle? Because this one is a "2 step" problem (not obvious) but to figure it out, they tell you the angle is 135º that means you should think: how far around the circle is that ? 135/360 which you could simplify to ⅜ next you notice the distance of the small part of the circle is 9 cm you should think: ⅜ of the circle = 9 cm or, better: ⅜ of circumference = 9 cm we could write that as ⅜ * C = 9 step 2: formula for circumference is C= 2 pi r if we replace C with 2 pi r in the other equation we get ⅜ * 2 * pi * r = 9 but you need to know a little algebra to "solve for r"

OpenStudy (gudden):

:D